OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3

OpenStudy (anonymous):

a=3--> a^2 =?

OpenStudy (anonymous):

-3^2

OpenStudy (anonymous):

ok, you skip the first step, but it's ok, -a^2= -3^2 =??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

-9, not 9